In 4 years, Shana`s mother will be twice shana`s age today. If shana will be 30 years old then, how old is shana`s mother now?
Answer by ankor@dixie-net.com(22740)

s = Shanna's age
m = mother's age
:
Write an equation for each statement
:
In 4 years, Shana`s mother will be twice shana`s age today.
m + 4 = 2s
If shana will be 30 years old then,
s = 30
how old is shana`s mother now?
m + 4 = 2(30)
m = 60 - 4
m = 56 is mothers age now